A sample of 20 observations is selected from normal population distribution with

population standard deviation 9 and sample mean 51. Determine the highest

value of population mean with 95% confidence interval (z-1.96). Round the

answer to the two decimal places.


1
Expert's answer
2022-05-05T06:47:59-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(51-1.96\times\dfrac{9}{\sqrt{20}}, 51+1.96\times\dfrac{9}{\sqrt{20}})"

"=(47.06, 54.94)"

The highest value of population mean with 95% confidence interval is "54.94."